# The Spacetime of Diffusion Models: An Information Geometry Perspective

### v2

@@ Appendix E Qualitative example of DiffED
Appendix F Experimental details
F.1 Toy Gaussian mixture
@@ F.3 Molecular data
In Section˜6.3, we want to interpolate between two low-energy states: ${\bm{x}}_{0}^{1}=(-2.55,2.7)$ and ${\bm{x}}_{0}^{2}=(0.95,-0.4)$. To avoid instabilities for $t\approx 0$, we represent them on the spacetime manifold as ${\bm{z}}_{1}=(-2.55,2.7,t_{\mathrm{min}})$, and ${\bm{z}}_{2}=(0.95,-0.4,t_{\mathrm{min}})$, where $\log\mathrm{SNR}(t_{\mathrm{min}})=7$. We then approximate the geodesic between them, by parametrizing $\bm{\gamma}$} as a cubic spline with 10 nodes and fixed endpoints ${\bm{\gamma}}_{0}={\bm{z}}_{1}$, and ${\bm{\gamma}}_{1}={\bm{z}}_{2}$ and discretize it into 512 points. We then optimize it by minimizing Proposition˜5.1 with the Adam optimizer with learning rate $\eta=0.1$ and take 10000 optimization steps, which takes roughly 6 minutes on an M1 CPU.
@@
To generate transition paths, we use Annealed Langevin dynamics (Algorithm˜1) with the geodesic discretized into $N=512$ points, $K=128$ Langevin steps for each point on the geodesic $\bm{\gamma}$}, and use $dt=0.0001$, i.e., requiring 65536 evaluations of the gradient of the auxiliary energy function. We generate 8 independent paths in parallel, which takes roughly 27 seconds on an M1 CPU.
@@
Appendix G Note on transition path sampling baselines
@@
Appendix H Expectation parameter estimation code

### v3

@@ Appendix F Qualitative example of DiffED
Appendix G Experimental details
G.1 Toy Gaussian mixture
@@ G.3 Molecular data
In Section˜6.3, we want to interpolate between two low-energy states: ${\bm{x}}_{0}^{1}=(-2.55,2.7)$ and ${\bm{x}}_{0}^{2}=(0.95,-0.4)$. To avoid instabilities for $t\approx 0$, we represent them on the spacetime manifold as ${\bm{z}}_{1}=(-2.55,2.7,t_{\mathrm{min}})$, and ${\bm{z}}_{2}=(0.95,-0.4,t_{\mathrm{min}})$, where $\log\mathrm{SNR}(t_{\mathrm{min}})=4$. We then approximate the geodesic between them by parametrizing $\bm{\gamma}$} as a cubic spline with 10 nodes and fixed endpoints ${\bm{\gamma}}_{0}={\bm{z}}_{1}$, and ${\bm{\gamma}}_{1}={\bm{z}}_{2}$ and discretize it into 128 points. We then optimize it by minimizing Proposition˜5.1 with the Adam optimizer with learning rate $\eta=0.1$ and take 10000 optimization steps, which takes roughly 4 minutes on an M1 CPU.
@@
To generate transition paths, we use Annealed Langevin dynamics (Algorithm˜1) with the geodesic discretized into $N=128$ points, $K=128$ Langevin steps for each point on the geodesic $\bm{\gamma}$}, and use $dt=0.0004$, i.e., requiring 16385 evaluations of the gradient of the auxiliary energy function per path. Generating 1000 independent paths in parallel takes 32 seconds on an M1 CPU and requires a total of 16,385,000 energy function evaluations.
@@
Appendix H Note on transition path sampling baselines
@@
Appendix I Expectation parameter estimation code